OpenStudy (anonymous):

Can someone show me how to factor 5(a+b)^2 - 20

OpenStudy (anonymous):

5(a+b)^2 - 20 =0 5(a+b)^2 = 20 (a+b)^2 =4 (a+b)^2 =2^2 a+b = +2 or -2

OpenStudy (watchmath):

\(5(a+b)^2-20\) \(5((a+b)^2-4\) \(5(a+b+2)(a+b-2)\)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Watchmath,how did you get 5(a+b)^2 -4